OpenStudy (anonymous):

simplify. with this one i got stuck halfway 2 /3a * 2/a^2 1) 2a/ 3a^2 * 6a/ 3a^2 2) 12a / 3a^2

OpenStudy (rsadhvika):

\(\large \frac{2}{3a} * \frac{2}{a^2}\)

OpenStudy (rsadhvika):

its like above ?

OpenStudy (anonymous):

yes

OpenStudy (rsadhvika):

ok, what are 1) and 2) you listed there

OpenStudy (rsadhvika):

are they options ?

OpenStudy (anonymous):

my steps

OpenStudy (rsadhvika):

ohk..

OpenStudy (anonymous):

but i think theyre wrong

OpenStudy (rsadhvika):

its a multiplication sign in between right ?

OpenStudy (anonymous):

yup

OpenStudy (rsadhvika):

its this, 2 /3a * 2/a^2 and not this 2 /3a + 2/a^2

OpenStudy (anonymous):

yup

OpenStudy (rsadhvika):

ok then easy, we just multiply numerators separately. then multiply denominators separately.

OpenStudy (rsadhvika):

\(\large \frac{2}{3a} * \frac{2}{a^2} \) step 1) \(\large \frac{2*2}{3a*a^2} \) step 2) \(\large \frac{4}{3a^3} \)

OpenStudy (rsadhvika):

we're done

OpenStudy (anonymous):

thats all? ok

OpenStudy (rsadhvika):

yes !